scott currently saves 15% of his income his annual income is 20925.00 how much more money would he be saving per month if he put aside 23% of his income instead ?
1
Expert's answer
2012-06-19T10:33:50-0400
The savings per month:
S1 = 20925.00 * 0.15 / 12 = 261.5625
And if he put
aside 23% of his income instead:
S2 = 20925.00 * 0.23 / 12 =
401.0625

S2 - S1 = 139.5000